OpenStudy (theopenstudyowl):

This will be the open thread for the SAT math for tonight! Super excited to work together with other people. I am going to post 10 different practice problems and one super hard problem that are all SAT math questions. Well go over answers to and tricks to solve them. Starts in 5 minutes...

OpenStudy (theopenstudyowl):

10 minutes actually

OpenStudy (theopenstudyowl):

lol

OpenStudy (haichi):

XD

OpenStudy (theopenstudyowl):

5 Minutes

OpenStudy (haichi):

*jeopardy theme song plays*

OpenStudy (theopenstudyowl):

hah

OpenStudy (theopenstudyowl):

Ok guys hopefully we can recruit more people in the future... I am going to hold these sessions at 10 every night pacifica time, but tonight is the first... I will hold them up until the March SAT and I am going to alternate subjects as well!

OpenStudy (theopenstudyowl):

So, I am going to post the first problem, just give me a second... We ill start basic and move up gradually.

OpenStudy (haichi):

ok kewl

OpenStudy (theopenstudyowl):

If the average of x, 5x and 6x is 8, what is the value of x?

OpenStudy (theopenstudyowl):

Try solving it without any answer choices first...

OpenStudy (haichi):

x=2?

OpenStudy (theopenstudyowl):

a. 1 b. 2 c. 3 d. 4 e. 5

OpenStudy (theopenstudyowl):

ok, let the other people answer if they want to for a minute and then well go over the correct answer

OpenStudy (haichi):

ok sorry :(

OpenStudy (theopenstudyowl):

its cool

OpenStudy (theopenstudyowl):

well if anyone else isnt going to answer, the correct answer is 2

OpenStudy (theopenstudyowl):

Haichi, u can share how u got the answer if u want to

OpenStudy (haichi):

I will

OpenStudy (theopenstudyowl):

cool!

OpenStudy (haichi):

so to get the answer it would be x+5x+6x=12x But there are three numbers: x, 5x and 6x so it would be \[12x \over 3\] and that would equal 4x

OpenStudy (haichi):

The average is 8 so \[8\over 4x \]=2

OpenStudy (theopenstudyowl):

yup perfect! If anyone has any questions about how he came to the answer he did, we can all go over it together..

OpenStudy (theopenstudyowl):

Ok, I guess we r on to the next question...

OpenStudy (theopenstudyowl):

If 3x^2=4y=12, what is the value of x^2y? a. 48 b. 36 c. 24 d. 12 e. 6

OpenStudy (haichi):

Should I say the answer?

OpenStudy (haichi):

With the explaination?

OpenStudy (theopenstudyowl):

yes

OpenStudy (haichi):

4y=12 so y= 3; now find value of x so 3x^2=12 then x^2=4 so x=2 if x=2 and y=3 2^(2)(3) or 2^6 This equals 64 so x^2y=64

OpenStudy (haichi):

I work it out on word so I dont cheat and look it up

OpenStudy (theopenstudyowl):

ok cool

OpenStudy (theopenstudyowl):

there is no worries about cheating hear man because were just practicing

OpenStudy (haichi):

ik

OpenStudy (theopenstudyowl):

feel free to look up formulas and stuff too

OpenStudy (haichi):

Just to clear things up

OpenStudy (theopenstudyowl):

ok one sec, give me a minute to type up the next question

OpenStudy (theopenstudyowl):

If (2x-5)(2x+5)=5, what is the value of 4x^2?

OpenStudy (theopenstudyowl):

-30 -20 10 20 30

OpenStudy (theopenstudyowl):

one of those are the answer...

OpenStudy (theopenstudyowl):

ok sorry guys

OpenStudy (theopenstudyowl):

bad connection

OpenStudy (haichi):

np

OpenStudy (theopenstudyowl):

this is going to be it for tonight because my connections sucks

OpenStudy (haichi):

Aww man :( I enjoyed this enrichment

OpenStudy (haichi):

(2x - 5)(2x + 5) = 5 4x^2 - 25 = 5 4x^2 = 30 x^2 = 15/2 x = +- (square root of 7.5) 4x = 4(square root of 7.5) or -4(square root of 7.5)

OpenStudy (theopenstudyowl):

ok tomorrow at 10 again

OpenStudy (haichi):

Awsome!

OpenStudy (theopenstudyowl):

alright see yall

OpenStudy (haichi):

buh bai nite sweet dreams don't let the little voids bite xD-from King.Void

OpenStudy (theopenstudyowl):

Happy studying